OpenStudy (banjoman):

Two angles of a triangle have the same measure and the third one is 33 degrees greater than the measure of each of the other two. Find the measure of the LARGEST angle in the triangle. Medal for answer and help

OpenStudy (johnweldon1993):

Hint* What do the interior angle of a triangle add up to?

OpenStudy (banjoman):

they add up to 66 I believe

OpenStudy (jackthegreatest):

let 2x = the combined measure of the two angles let x + 33=the measure of the third angle ...180 degrees in a triangle... 2x + (x+3) = 180

OpenStudy (jackthegreatest):

there is your equation, can you solve?

OpenStudy (banjoman):

i got 59?

OpenStudy (jackthegreatest):

whoops i had a typo

OpenStudy (jackthegreatest):

i meant 2x + (x + 33) = 180

OpenStudy (jackthegreatest):

so sorry, x anyways x = 49

OpenStudy (jackthegreatest):

add 33 to 49, and that is your answer